2. The Motor Transport Authority (MTA) expects that new car sales will decrease from 48,500 cars in December to 46,950 cars in January due to 5% Value Added Tax (VAT) in 2018, which will be completely passed on to consumers. Calculate PED, state whether it is elastic or inelastic, and, state what manufacturers and dealers should do to maximize revenue.


1
Expert's answer
2021-04-23T12:07:09-0400

Solution:

Price Elasticity of Demand(PED) = "\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


% change in qty demanded = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \times 100 = \frac{46950 -48500}{(46950+48500)/2 } \times 100"


"\frac{-1550}{47725} \times 100 = -0.03248\times 100 = -3.248\%"


% change in price = 5%



Price Elasticity of Demand(PED) = "\frac{-3.248\%}{5\%} = -0.65"


PED = 0.65


The elasticity of demand is less than 1, which means that PED is inelastic and a normal good.

 

Since PED is inelastic, this means that a price increase will result in a smaller percentage decrease in the number of cars sold. Therefore, manufacturers and dealers should raise the price of their cars which will ultimately increase the total revenue.